OpenStudy (rayford):

What is the sum of a 28–term arithmetic sequence where the first term is 12 and the last term is 255?

OpenStudy (rayford):

wait..but i don't understand the formula.

hartnn (hartnn):

for more on this topic: Tutorial on AP,GP,HP. http://openstudy.com/study#/updates/503bb2a0e4b007f9003103b0

OpenStudy (rayford):

couldn't you just work it out? i learn best from rereading.

hartnn (hartnn):

sure , sum of n terms = n/2(1st term + last term)

hartnn (hartnn):

= 28/2 (12+255)

hartnn (hartnn):

=14*(267) =3738 got it ?

OpenStudy (rayford):

wow i actually do..

OpenStudy (rayford):

thank you so much..

hartnn (hartnn):

thats great :)

hartnn (hartnn):

welcome ^_^

OpenStudy (rayford):

wait..

OpenStudy (rayford):

for the future..how do i tag people in a problem if noone will answer?

hartnn (hartnn):

`@hartnn`

hartnn (hartnn):

@rayford

hartnn (hartnn):

tag me now

OpenStudy (rayford):

ok, i hate to be a burden but..for the next 2 days is it ok..if you be my tutor?

hartnn (hartnn):

if i m online, i will be....else there are many to help u here :)

OpenStudy (rayford):

ok..but. do you think you will be online at 8 tomorrow?

hartnn (hartnn):

my time zone maybe different, its 8 am here now....

OpenStudy (rayford):

oh man..it's 10:21 pm here

hartnn (hartnn):

so means 10 hrs from now.... means at 6pm for me....most probably, i will be online, but cannot guarantee anything....

OpenStudy (rayford):

i absolutely understand.